Inside a box are a R10 note, a R20 note, a R50 note and a R100 note. Consider the experiment in which two notes are drawn from the box without replacement. Let đť‘‹ be the random variable that gives the value of the smaller (Rands) note. Construct the probability mass function for đť‘‹


1
Expert's answer
2021-05-19T17:34:26-0400

The total outcomes of drawing two notes from the box =6

They are -{R10,R20},{R10,R50},{R10,R100},{R20,R50},{R20,R100},{R50,R100}


When outcomes are -{R10,R20},{R10,R50},{R10,R100}


"X=10, \text{ and } P(X=10)=\dfrac{3}{6}=0.5"


When outcomes are -{R20,R50},{R20,R100}


"X=20 \text{ and }P(X=20)=\dfrac{2}{6}=0.33"


When outcome is {R50,R100}


"X=50 \text{ and }P(X=50)=\dfrac{1}{6}=0.166"


Therefore the probability mass function for X is as follows: